A multiscale method for separation processes in chemical engineering

Karsten Urban RWTH Aachen, Germany

Abstract

Many common examples of fluid mixtures encountered in chemical and hydrocarbon processing industries contain far too many components as possible for a detailed compositional analysis and subsequent modelling. However, the knowledge of the particular amount of these components plays an important role in separation processes as e.ġ. in a single equilibrium flash unit. These processes are described in terms of a time dependent system of differential algebraic equations.

In this talk we introduce a multiscale method for such a problem, which aims at significantly reducing the size of the underlying discrete systems. Moreover, the error caused by compression can be controlled and the multiscale representation of the data ensures that at any stage the results can be expressed in terms of all the original components. We indicate the analysis of the algorithm and present some preliminary numerical results.

The presented results are joint work with Wolfgang Dahmen, Wolfgang Marquardt and Rüdiger von Watzdorf.